I have a couple questions that are hard for me to answer.
1. what is the answer to this? log(2)(x^2+2x-44)=2?

Answer by stanbon(48569)

log(2)(x^2+2x-44)=2?
That says x^2+2x-44 = 2^2
x^2+2x-48 = 0
(x+8)(x-6) = 0
x = -8 ; x = 6
-----------
Checking:
x = -8
log(2)(64-16-44) = 2
log(2)(4) = 2
4 = 2^2
-----------------
Checking:
x = 6
log(2)(36+2*6-44) = 2
log(2)(4) = 2
4 = 2^2
--------------
Cheers,
Stan H.